DISCRETE MATH

If the public key e = 11, and p = 3, q = 5. What i
Answer #1

p=3

q=5

n=p*q = 3*5 = 15

\phi(n) = (p-1) * (q-1)= 2 * 4 = 8

e=11

choosedsuchthat , d*e\equiv 1(\mod \phi(n))

d = 3i.e. ,3 * 11 = 33 \equiv 1 \mod 8

d= 3